OpenStudy (anonymous):

Which of the following is a solution to cos(2x) = 1/2 60° 120° 150° 330°

OpenStudy (anonymous):

@e.mccormick

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@iGreen

OpenStudy (jdoe0001):

\(\bf cos(2x)=\cfrac{1}{2}\implies cos^{-1}[cos(2x)]=cos^{-1}\left( \cfrac{1}{2} \right)\implies 2x=cos^{-1}\left( \cfrac{1}{2} \right) \\ \quad \\ x=\cfrac{cos^{-1}\left( \frac{1}{2} \right)}{2}\)

OpenStudy (anonymous):

I cant find the degrees

OpenStudy (jdoe0001):

well... what's the angle that \(\bf cos^{-1}\left( \cfrac{1}{2} \right)\)?

OpenStudy (jdoe0001):

is positive, and cosine is only positive in the 1st and 4th quadrants

OpenStudy (anonymous):

150 degrees

OpenStudy (jdoe0001):

150 is on the 2nd quadrant

OpenStudy (anonymous):

so its either a or d

OpenStudy (jdoe0001):

in case you don't have a Unit Circle -> http://www.sciencedigest.org/UnitCircle.gif

OpenStudy (anonymous):

how is 150 the second quadrant?

OpenStudy (anonymous):

maybe im looking at it wrong

OpenStudy (jdoe0001):

150 = 90 + 60 

OpenStudy (anonymous):

so 60 is in the 1st quadrant?

OpenStudy (anonymous):

60 degrees

OpenStudy (jdoe0001):

yeap.. notice your Unit Circle 

OpenStudy (anonymous):

so the answer is in the 4th quadrant?

OpenStudy (jdoe0001):

either of those have a cosine of 1/2 thus let's say, using \(\pm 60\) in this case \(\bf cos(2x)=\cfrac{1}{2}\implies cos^{-1}[cos(2x)]=cos^{-1}\left( \cfrac{1}{2} \right)\implies 2x=cos^{-1}\left( \cfrac{1}{2} \right) \\ \quad \\ x=\cfrac{cos^{-1}\left( \frac{1}{2} \right)}{2}\implies x=\cfrac{\cancel{\pm60}}{\cancel{2}}\to \pm30^o\)

OpenStudy (anonymous):

So it is 60 degrees thanks